You read in a newspaper that the nominal interest rate is 12 percent per year in Canada and 8 percent per year in the United States. Suppose that international capital flows equalize the real interest rates in the two countries and that purchasing power parity holds.

a. Using the Fisher equation, what can you infer about expected inflation in Canada and in the United States?

b. What can you infer about the expected change in the exchange rate between the Canadian dollar and the U. S. dollar?

c. A friend proposes a get-rich-quick scheme: borrow from a U. S. bank at 8 percent, deposit the money in a Canadian bank at 12 percent, and make a 4 percent profit. What’s wrong with this scheme?
Answer #1

a.

Using Fisher's equation,

nominal interest rate ~ real interest rate + inflation.

real interest rate = nominal interest rate - inflation

Thus, as the real interest rate in US and Canada are same,

Nominal rate Canada - Inflation Canada = Nominal Rate US - Inflation US

12% - inflation Canada = 8% - Inflation US

Inflation Canada = 4% + Inflation US

Thus inflation in Canada will be 4% higher than in US

b. As inflation in Canada is higher, its value will depreciate again the US dollar. i.e the US dollar will get stronger and Canadian dollar will get weak. The change in the value will be 4%.

c.

Let 1 USD = 1 CAD and amount borrowed be $1 in US for a year.

Amount payable in US after 1 year = $1 * (1 + 8%) = $1.08

Amount converted to CAD = CAD 1

Amount in CAD after 1 year = CAD 1*(1+12%) = CAD 1.12

New Exchange rate after 1 year = 1 * (1 + 4%) = 1.04 i.e 1 USD = 1.04 CAD (i.e USD costs more in CAD)

Amount in USD converted from CAD = 1.12/1.04 = 1.0769 ~ 1.077 USD

Thus, this scheme actually loses money as the amount of loan to be returned is 1.08 USD which is more than 1.077USD earned from CAD interest rates. (This is even without considering the exchange rate transaction fees etc.)